OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 2y = 6 x + 3y = −1 −2 2 5 −5 11. [6.03] Solve the following system. x − y = 10 x + y = 8 (−9, 1) (9, −1) (9, 1) (−1, 9)

OpenStudy (michele_laino):

question #1 here I solve the second equation for x, and I get: x=-1-3y then I substitutte into the first equation, and I get: \[2\left( { - 1 - 3y} \right) + 2y = 6\] what is y?

OpenStudy (triciaal):

2x + 2y = 6 x + 3y = −1 can simplify the first eqn by dividing by 2 so x + y = 3 subtract from the 2nd eqn 2y = -4 so y = -2 substitute in either original x = -1 -3(-2) = 5 check solution

OpenStudy (triciaal):

x − y = 10 x + y = 8 coefficient of y already equal and opposite so just add eqns to eliminate y 2x = 18 so x = 9 substitute for x in either y = -1